OpenStudy (anonymous):

Graph the following linear inequality: 6 x + 7 y < 7

OpenStudy (anonymous):

my plots: x=7/6, y=-1

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

Very close :) 0 + 7y = 7 /7 /7 y = 1

OpenStudy (anonymous):

So from (0,1), I got up 6 & over 1?

jigglypuff314 (jigglypuff314):

hmmm... solve for y 7y = -6x + 7 y = -6/7 x + 1 so slope = - 6/7 so from (0, 1) down 7 over 6

OpenStudy (anonymous):

ooooooh ok! Next one is the last one & it has a fraction! :O

OpenStudy (anonymous):

I says it's wrong, I put it on the negative side & everything :(

jigglypuff314 (jigglypuff314):

hmmm...?

OpenStudy (anonymous):

Am I suppose to shade it at the top?

jigglypuff314 (jigglypuff314):

below... :/